# The Optimization Landscape of SGD Across the Feature Learning Strength

### v1

@@ 1.1 Related Works
Chizat et al. (2019) introduced the laziness parameter $\alpha$ in their seminal work. Mei et al. (2018); Rotskoff & Vanden-Eijnden (2018) introduced networks in mean-field parameterization. Geiger et al. (2020) introduced the correct scaling of $\alpha$ with $N$ as $\alpha=(\gamma\sqrt{N})^{-1}$ to achieve feature learning at infinite width. They also highlighted that $\gamma$ can be identified as the parameter that controls feature learning. Yang & Hu (2021) expanded this idea to deep networks of arbitrary architecture and coined the term $\mu$P. Bordelon & Pehlevan (2022) described this limit in terms of dynamical mean field theory. Extensions to infinite depth were performed in Bordelon et al. (2023) and Yang et al. (2023b) and to infinite attention in Bordelon et al. (2024c). We give a review of network parameterizations in Appendix B. A key motivation for this work is that $\gamma$ is the main free parameter in the DMFT description of infinite-width feature learning networks, warranting a more thorough empirical investigation.
@@ 2 Setup and Notation
We train neural networks $f(\bm{x};\bm{\theta})$ parameterized using $\mu$P scaling, with the output further downscaled by a factor of $\gamma$ [^2] We verify our $\mu$P implementation, confirming consistency of dynamics across width in Appendix F. We denote the width of a given network by $N$ and the depth of a feed-forward network by $L$ [^3] Here, depth counts the number of weight matrices, e.g. $f(\mathbf{x})=W_{2}\sigma(W_{1}\mathbf{x})$ has depth $L=2$. We train online on batches $\mathcal{B}_{t}=\{\bm{x}_{\mu},\bm{y}_{\mu}\}_{\mu=1}^{B}$ batch with size $B$ for $t\in\{1,\dots,T\}$ time steps. The loss of each batch is given by a mean of per-example losses:
@@
We will consider both mean-squared-error and cross-entropy loss for $\ell$. In this work we focus on optimizing the parameters by vanilla SGD, as even in this simple setting, the effect of $\gamma$ is not well-understood. For small batch sizes, the the SGD effects are controlled by the ratio $\eta/B$ Jastrzebski et al. (2017); Smith et al. (2020); Sclocchi et al. (2023). We verify this this in Appendix H.
@@ 3.1 Phase Portrait of $\eta$ with $\gamma$
Figure 2: a) Online loss curves of CNN trained on CIFAR-5M across $\gamma$, at the second largest convergent $\eta$. $\eta$ scales as $\gamma^{2}$ for $\gamma<1$ and $\gamma^{2/L}$ for $\gamma>1$. We also observe that very large $\gamma$ have a long period of flat loss before a drop. We overlay dashed lines to highlight the different power law scalings of loss with training time observed in the lazy and rich regime. b) Final test accuracy as a function of $\gamma$. We see that, as long as one trains long enough, larger $\gamma$ yields equal or better generalization to $\gamma=1$, but the returns are marginal past some point. We verify this for other networks in Appendix I.3. Error bars show the small variance over three initializations.
@@
By centering, near initialization $\ell^{\prime}$ is $\Theta_{\gamma}(1)$. The left hand side is thus $\Theta(\gamma^{-2})$, so the minimum learning rate to see the loss drop in $T$ steps scales as $\eta\sim\gamma^{2}/T$. We see this on the left region of the plots in Figure 1 and 1. We verify that networks of different but sufficiently small $\gamma$ have the same loss dynamics in Appendix I.1, indicating that the lazy limit is reached by realistic networks.
@@ 3.2 Hessian Scaling
Figure 3: a) The top 20 eigenvalues of the Hessian at the end of training vs $\gamma$ for an NN trained with MSE. We clearly see two regimes. For small $\gamma$ we see that $\lambda_{\max}$ b) The eigenvalue vs time across several value of $\gamma$. c) The top eigenvalues across time for a lazy network. In the lazy setting, we see ten outliers, equal to the number of classes. The Hessian otherwise does not change in MSE. d) The same for a rich network. At late time, rather than seeing a small set of outliers, we see many eigenvalues grow to a sizeable range. Further plots are given in Appendix J.
@@ 3.3 Dynamical Phenomena
Figure 4: a) At sufficiently small $\gamma$, we observe that the loss catapults. Small $\gamma$ incur a larger catapult. Further catapult plots are in Appendix K. b) At large $\gamma$, we do not observe catapults. At the optimal learning rate, the loss stays near a saddle for an extended period of time, before suddenly dropping. We show that during this period, the alignment between the final-layer kernel and the task grows before the loss drops. Kernel-target alignment is defined in Appendix L. Further silent alignment plots are in Appendix M. We also see characteristic step-wise loss drops in this regime.
@@
At very large $\gamma$, we often observe loss trajectories following a characteristic staircase pattern, with one or more plateaus punctuated by sudden drops. This is a prediction of linear network theory Saxe et al. (2014); Simon et al. (2023a) that seems to hold even for very realistic models. We see these step-wise trajectories for MLPS in Figures 4, 27 and 35, CNNs in Figures 27 and 34 and ViTs in Figure 12. We show that, upon rescaling time as $\tau=\eta t/\gamma$, the early time dynamics coincide. The first loss drop goes as $t\sim\gamma$. We explain the origin of this timescale in Section 4.
@@ 3.4 Comparing Learned Functions and Representations
Figure 5: a) Function outputs between pairs of networks, as described in the main text. We see that lazy networks have nearly identical function outputs, and that rich networks agree in their function outputs at the end of training as well. See Appendix N for further plots. b) We study the kernel-target alignment of final-layer representations across $\gamma$ at the end of training.
@@ 4.1 Scalings for MSE loss
Table 1: Learning rate scalings for our toy models for two loss functions in the lazy and ultra-rich regimes. All values come out of our one-parameter model except $\eta_{\text{crit}}$ for MSE loss, which comes from the two-parameter model given in Appendix D. Predicted scalings match NN experiments.
@@
This phase ends when the function value $\frac{w^{L}}{\gamma}$ reaches order unity. In order for this to occur, we need $\eta\sim T^{-1}\gamma\log\gamma$ if $L=2$ and $\eta\sim\gamma/T$ if $L\geq 3$. Neglecting the logarithmic factor, we find a lower bound $\eta_{\text{min}}\sim\frac{\gamma}{T}$. Note as a consequence that the early-time dynamics are approximately $\gamma$-invariant after the rescaling $\tau\equiv\eta t/\gamma$. We verify this in Figure 4, 5, and Appendix I.2.
@@
The catapult effect generally requires two parameter dimensions to describe Damian et al. (2023), so our one-parameter model cannot capture it. In Appendix K, we describe and simulate a minimal two-parameter model that exhibits catapults. We find that for MSE loss, catapults occur in a narrow band $\eta_{\text{crit}}<\eta<\eta_{\max}$, with $\eta_{\text{crit}}\sim\gamma^{2}$ just like $\eta_{\text{max}}$.
Figure 6: Simulations of our one-parameter model with MSE loss (left; see Equation 4) and cross-entropy loss (right; see Equation 9) with varying $\gamma,\eta$ recover our principal phase portrait. For simulations of a two-parameter model which exhibits true catapults, see Figures 8 and 9.
@@ 4.2 Scalings for Cross-Entropy Loss
[…]
We perform simulations of our one-parameter model in a fashion analogous to our parameter sweeps for deep networks for $T=10^{3}$ steps of gradient descent. The results, plotted in Figure 6, confirm our analytical scalings and match our phase portraits for deep networks (e.g. Figure 1).
@@ acknowledgements
The authors are thankful to Blake Bordelon, Daniel Kunin, Katherine Lee, Allan Raventos, Jascha Sohl-Dickstein, Jacob Zavatone-Veth, and Brian Zhang for helpful discussions. AA thanks Blake Bordelon and Jeremy Cohen for in-depth discussions that motivated many of the questions investigated in this paper. We thank Blake Bordelon and Jacob Zavatone-Veth for comments on a draft. AA was supported by a Yaser Abu-Mostafa Fellowship from the Fannie and John Hertz Foundation. JS gratefully acknowledges support from the National Science Foundation Graduate Fellow Research Program (NSF-GRFP) under grant DGE 1752814. CP is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, and a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Discussions contributing to this work took place at the Kavli Institute for Theoretical Physics’ 2023 program on “Deep Learning from the Perspective of Physics and Neuroscience.” AA and JS thank the builders of a particular house in Reston, VA which, to their surprise, they discovered they had both lived in at different times.
@@ Appendix A Experimental Details
A.1 Comments on Numerical Stability in the Lazy Limit
Here we detail a ubiquitous problem that arose in the training of lazy neural networks at small $\gamma$. To our knowledge, this problem has not been highlighted in the literature. At very small values of $\gamma$, approximately $10^{-3}$, the appropriate learning rate in the lazy limit scales as $\eta\sim\gamma^{2}$. This leads to weight updates on the order of $10^{-6}$ or less. Machine epsilon for float32 is near $10^{-7}$. Because the final output of the function involves the subtraction in equation 10, the weight updates matter in both relative and absolute magnitude. For this reason, we find that we needed to go to double precision to reliably obtain curves below $\gamma=10^{-3}$. In practice, we find that a value of $\gamma=10^{-3}$ or even $10^{-2}$ is essentially enough to reach the lazy limit in the online setting for essentially all tasks listed. See Appendix I.1 for a check of this.
@@ Appendix B Review of NTK Parameterization and $\mu$-Parameterization
For detailed discussion of $\mu$P more generally, see Yang & Hu (2021); Yang et al. (2021); Geiger et al. (2020); Bordelon & Pehlevan (2022). The aim of this section is conceptual clarity on the difference between NTK parameterization and $\mu$-parameterization, as well as to give some degree of motivation of the latter. See also the recent accessible works of Yang et al. (2023a) and Chizat & Netrapalli (2023).
@@ D.3 Simulations of the two-parameter toy model
Appendix E Additional ResNet and ViT Plots

### v3

@@ 1.1 Related Works
Chizat et al. (2019) introduced the laziness parameter $\alpha$ in their seminal work. Mei et al. (2018); Rotskoff & Vanden-Eijnden (2018) introduced networks in mean-field parameterization. Geiger et al. (2020) introduced the correct scaling of $\alpha$ with $N$ as $\alpha=(\gamma\sqrt{N})^{-1}$ to achieve feature learning at infinite width. They also highlighted that $\gamma$ can be identified as the parameter that controls feature learning. Yang & Hu (2021) expanded this idea to deep networks of arbitrary architecture and coined the term $\mu$P. Bordelon & Pehlevan (2022) described this limit in terms of dynamical mean field theory. Extensions to infinite depth were performed in Bordelon et al. (2023) and Yang & Littwin (2023) and to infinite attention in Bordelon et al. (2024c). We give a review of network parameterizations in Appendix B. A key motivation for this work is that $\gamma$ is the main free parameter in the DMFT description of infinite-width feature learning networks, warranting a more thorough empirical investigation.
@@ 2 Setup and Notation
We train neural networks $f(\bm{x};\bm{\theta})$ parameterized using $\mu$P scaling, with the output further downscaled by a factor of $\gamma$ [^2] We verify our $\mu$P implementation, confirming consistency of dynamics across width in Appendix I. We denote the width of a given network by $N$ and the depth of a feed-forward network by $L$ [^3] Here, depth counts the number of weight matrices, e.g. $f(\mathbf{x})=W_{2}\sigma(W_{1}\mathbf{x})$ has depth $L=2$. We train online on batches $\mathcal{B}_{t}=\{\bm{x}_{\mu},\bm{y}_{\mu}\}_{\mu=1}^{B}$ batch with size $B$ for $t\in\{1,\dots,T\}$ time steps. The loss of each batch is given by a mean of per-example losses:
@@
We will consider both mean-squared-error and cross-entropy loss for $\ell$. In this work we focus on optimizing the parameters by vanilla SGD, as even in this simple setting, the effect of $\gamma$ is not well-understood. For small batch sizes, the the SGD effects are controlled by the ratio $\eta/B$ Jastrzebski et al. (2017); Smith et al. (2020); Sclocchi et al. (2023). We verify this this in Appendix K.
@@ 3.1 Phase Portrait of $\eta$ with $\gamma$
Figure 2: a) Online loss curves of CNN trained on CIFAR-5M across $\gamma$, at the second largest convergent $\eta$. $\eta$ scales as $\gamma^{2}$ for $\gamma<1$ and $\gamma^{2/L}$ for $\gamma>1$. We also observe that very large $\gamma$ have a long period of flat loss before a drop. We overlay dashed lines to highlight the different power law scalings of loss with training time observed in the lazy and rich regime. b) Final test accuracy as a function of $\gamma$. We see that, as long as one trains long enough, larger $\gamma$ yields equal or better generalization to $\gamma=1$, but the returns are marginal past some point. We verify this for other networks in Appendix L.3. Error bars show the small variance over three initializations.
@@
By centering, near initialization $\ell^{\prime}$ is $\Theta_{\gamma}(1)$. The left hand side is thus $\Theta(\gamma^{-2})$, so the minimum learning rate to see the loss drop in $T$ steps scales as $\eta\sim\gamma^{2}/T$. We see this on the left region of the plots in Figure 1 and 1. We verify that networks of different but sufficiently small $\gamma$ have the same loss dynamics in Appendix L.1, indicating that the lazy limit is reached by realistic networks.
@@ 3.2 Hessian Scaling
Figure 3: a) The top 20 eigenvalues of the Hessian at the end of training vs $\gamma$ for an NN trained with MSE. We clearly see two regimes. For small $\gamma$ we see that $\lambda_{\max}$ b) The eigenvalue vs time across several value of $\gamma$. c) The top eigenvalues across time for a lazy network. In the lazy setting, we see ten outliers, equal to the number of classes. The Hessian otherwise does not change in MSE. d) The same for a rich network. At late time, rather than seeing a small set of outliers, we see many eigenvalues grow to a sizeable range. Further plots are given in Appendix M.
@@ 3.3 Dynamical Phenomena
Figure 4: a) At sufficiently small $\gamma$, we observe that the loss catapults. Small $\gamma$ incur a larger catapult. Further catapult plots are in Appendix N. b) At large $\gamma$, we do not observe catapults. At the optimal learning rate, the loss stays near a saddle for an extended period of time, before suddenly dropping. We show that during this period, the alignment between the final-layer kernel and the task grows before the loss drops. Kernel-target alignment is defined in Appendix P. Further silent alignment plots are in Appendix Q. We also see characteristic step-wise loss drops in this regime.
@@
At very large $\gamma$, we often observe loss trajectories following a characteristic staircase pattern, with one or more plateaus punctuated by sudden drops. This is a prediction of linear network theory Saxe et al. (2014); Simon et al. (2023a) that seems to hold even for very realistic models. We see these step-wise trajectories for MLPS in Figures 4, 31 and 41, CNNs in Figures 31 and 40 and ViTs in Figure 16. We show that, upon rescaling time as $\tau=\eta t/\gamma$, the early time dynamics coincide. The first loss drop goes as $t\sim\gamma$. We explain the origin of this timescale in Section 4.
@@ 3.4 Comparing Learned Functions and Representations
Figure 5: a) Function outputs between pairs of networks, as described in the main text. We see that lazy networks have nearly identical function outputs, and that rich networks agree in their function outputs at the end of training as well. See Appendix R for further plots. b) We study the kernel-target alignment of final-layer representations across $\gamma$ at the end of training.
@@ 4.1 Scalings for MSE loss
Table 1: SGD learning rate scalings for our toy models for two loss functions in the lazy and ultra-rich regimes. All values come out of our one-parameter model except $\eta_{\text{crit}}$ for MSE loss, which comes from the two-parameter model given in Appendix D. Predicted scalings match NN experiments.
@@
This phase ends when the function value $\frac{w^{L}}{\gamma}$ reaches order unity. In order for this to occur, we need $\eta\sim T^{-1}\gamma\log\gamma$ if $L=2$ and $\eta\sim\gamma/T$ if $L\geq 3$. Neglecting the logarithmic factor, we find a lower bound $\eta_{\text{min}}\sim\frac{\gamma}{T}$. Note as a consequence that the early-time dynamics are approximately $\gamma$-invariant after the rescaling $\tau\equiv\eta t/\gamma$. We verify this in Figure 4, 5, and Appendix L.2.
@@
The catapult effect generally requires two parameter dimensions to describe Damian et al. (2023), so our one-parameter model cannot capture it. In Appendix N, we describe and simulate a minimal two-parameter model that exhibits catapults. We find that for MSE loss, catapults occur in a narrow band $\eta_{\text{crit}}<\eta<\eta_{\max}$, with $\eta_{\text{crit}}\sim\gamma^{2}$ just like $\eta_{\text{max}}$.
Figure 6: Simulations of our one-parameter model with MSE loss (left; see Equation 4) and cross-entropy loss (right; see Equation 9) with varying $\gamma,\eta$ recover our principal phase portrait. Simulations run with $L=5$. For simulations of a two-parameter model which exhibits true catapults, see Figures 8 and 9.
@@ 4.2 Scalings for Cross-Entropy Loss
The main effect of this difference is seen in the lazy ($\gamma\ll 1$) regime when $\eta$ is too large for stability at the minimizer. Once $w-1$ has grown to order just larger than $\gamma$ (and we are still seeing linearized dynamics), it begins to experience constant-sized restoring steps $\delta w\sim-\frac{\eta L}{\gamma}\operatorname{sign}(w-1)$. [^5] Contrast this with the linear-sized restoring steps $\delta w\sim-\frac{L^{2}}{\gamma^{2}}\delta w$ which one sees with MSE. It is only if $\eta\sim\gamma$ or larger that this step is large enough to lead to a $\Theta_{\gamma}(1)$ change in $w$, upon which we escape our initial linear region and can diverge. The result is that the maximum stable learning rate for our toy model in the lazy regime is $\eta_{\text{crit}}\sim\gamma^{2}$, but the maximum nondivergent learning rate is $\eta_{\text{max}}\sim\gamma$. In our minimal two-parameter model described in Appendix D, we find that the loss does eventually settle back down and converge for $\eta_{\text{crit}}<\eta<\eta_{\max}$, and that this triangle of the phase diagram is there a true catapult regime. All the above scalings are summarized in Table 1.
We perform simulations of our one-parameter model in a fashion analogous to our parameter sweeps for deep networks for $T=10^{3}$ steps of gradient descent. The results, plotted in Figure 6, confirm our analytical scalings and match our phase portraits for deep networks (e.g. Figure 1). Finally, we also provide a similar analysis in the case of the Adam optimizer in Appendix Appendix F.
@@ acknowledgements
The authors are thankful to Blake Bordelon, Daniel Kunin, Katherine Lee, Allan Raventos, Jascha Sohl-Dickstein, Jacob Zavatone-Veth, and Brian Zhang for helpful discussions. AA thanks Blake Bordelon and Jeremy Cohen for in-depth discussions that motivated many of the questions investigated in this paper. We thank Blake Bordelon and Jacob Zavatone-Veth for comments on a draft. AA was supported by a Yaser Abu-Mostafa Fellowship from the Fannie and John Hertz Foundation. AM acknowledges the support of a Kempner Institute Graduate Research Fellowship. JS gratefully acknowledges support from the National Science Foundation Graduate Fellow Research Program (NSF-GRFP) under grant DGE 1752814. C.P. is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, DARPA grant DIAL-FP-038, a Sloan Research Fellowship, and The William F. Milton Fund from Harvard University. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Discussions contributing to this work took place at the Kavli Institute for Theoretical Physics’ 2023 program on “Deep Learning from the Perspective of Physics and Neuroscience.” AA and JS thank the builders of a particular house in Reston, VA which, to their surprise, they discovered they had both lived in at different times.
@@ Appendix A Experimental Details
A.1 Details on the Parameterization
On top of the centering operation explained above, ee parameterize all our networks under $\mu$P (and Depth-$\mu$P, respectively) parameterization. More concretely, we define our $L$ layer neural network of width $N$ $f(x)$ with nonlinearity $\phi(x)$ as:
$h^{1}(x)$ $=\frac{1}{\sqrt{D}}W^{0}x$ (11)
$h^{\ell+1}(x)$ $=\tau h^{\ell}(x)+\frac{1}{\sqrt{N}L^{\alpha}}W^{\ell}\phi(h^{\ell}(x))$ (12)
$f(x)$ $=\frac{1}{\gamma N}W^{L}\phi(h^{L}(x))$ (13)
where $D$ is the input dimension, and for $\tau=0$ and $\alpha=0$ we recover $\mu$P, and for $\tau=1$ and $\alpha=\frac{1}{2}$ we recover Depth-$\mu$(Bordelon et al., 2023; Yang & Littwin, 2023). When running SGD, we update the weights with a global learning rate going as $\eta N$. That is:
$W^{\ell}_{t+1}=W^{\ell}_{t}-N\eta\nabla\mathcal{L}$ (14)
Similar modifications exist for Adam, see Yang & Littwin (2023).
In the case of Transformers, we additionally change the attention temperature to $\frac{1}{d_{q}}$ following (Yang et al., 2022). We further detail and contrast this parameterization with the NTK parameterization in Appendix B.
A.2 Comments on Numerical Stability in the Lazy Limit
Here we detail a ubiquitous problem that arose in the training of lazy neural networks at small $\gamma$. To our knowledge, this problem has not been highlighted in the literature. At very small values of $\gamma$, approximately $10^{-3}$, the appropriate learning rate in the lazy limit scales as $\eta\sim\gamma^{2}$. This leads to weight updates on the order of $10^{-6}$ or less. Machine epsilon for float32 is near $10^{-7}$. Because the final output of the function involves the subtraction in equation 10, the weight updates matter in both relative and absolute magnitude. For this reason, we find that we needed to go to double precision to reliably obtain curves below $\gamma=10^{-3}$. In practice, we find that a value of $\gamma=10^{-3}$ or even $10^{-2}$ is essentially enough to reach the lazy limit in the online setting for essentially all tasks listed. See Appendix L.1 for a check of this.
@@ Appendix B Review of NTK Parameterization and $\mu$-Parameterization
For detailed discussion of $\mu$P more generally, see Yang & Hu (2021); Yang et al. (2021); Geiger et al. (2020); Bordelon & Pehlevan (2022). The aim of this section is conceptual clarity on the difference between NTK parameterization and $\mu$-parameterization, as well as to give some degree of motivation of the latter. See also the recent accessible works of Yang et al. (2023) and Chizat & Netrapalli (2023).
@@ D.3 Simulations of the two-parameter toy model
Appendix E Scaling Relationships are Unaffected by Learning Rate Warmup and Decay
In this section, we empirically verify that adding a warmup and cosine decay to the learning rate, as common in more modern deep learning practice, does not affect the key scaling relationships observed. The models are trained using learning rate warmup for approximately $5\%$ of the total number of steps, followed by a cosine decay until the end of training. All models are still trained with SGD. In the next section, we analyze the effect of alternative optimizers such as Adam and SignSGD.
Figure 10: Empirical runs of a 3 layer and readout ConvNet without residuals model trained with SGD with MSE loss (left; see Equation 4) and cross-entropy loss (right; see Equation 9) with varying $\gamma,\eta$ on CIFAR-5m.
Appendix F Scaling Relationships for Adam
All analysis thus far treats MSE loss. However, in practice, Adam is a more common choice of optimizer when training large models, so it is worth asking how our story changes when switching SGD to Adam. We repeat our analysis our toy model for SignSGD — which is equivalent to Adam without momentum — in Appendix F. We find a very similar story with different scaling exponents: $\eta_{\text{min}},\eta_{\text{max}}\sim\gamma$ when $\gamma\ll 1$ and $\eta_{\text{min}},\eta_{\text{max}}\sim\gamma^{1/L}$ when $\gamma\gg 1$. Remarkably, unlike with SGD, the upper and lower scaling exponents are the same in the ultra-rich regime ($\gamma\gg 1$) due to the fact that Adam escapes the saddle near initialization in $O(1)$ time. This suggests that Adam may be the optimizer of choice for future investigations into the ultra-rich regime.
In this appendix, we obtain both empirical and analytic phase plots and scaling relationships between $\eta$ and $\gamma$ under the Adam optimizer. We present a variant of the one-parameter model in Section 4 for training with SignSGD (Kingma, 2014). Adam can be thought of as essentially SignSGD with momentum. The predictions of the one-parameter model are then validated in via simulation.
F.1 Theoretical Phase Plot for SignSGD
Adam is essentially SignSGD (Bernstein et al., 2018) with momentum, and the effect of this momentum will only change our story up to order-unity prefactors, so we instead study SignSGD. The update rule for SignSGD is
$\theta\mapsto\theta-\eta\cdot\operatorname{sign}\left(\nabla_{\theta}\mathcal{L}\right),$ (23)
where $\theta$ is a vector of parameters and $\mathcal{L}$ is the loss. As in Section 4, we wish to determine the viable range of learning rates for training the centered, rescaled target function $\tilde{f}(w)=\gamma^{-1}(w^{L}-1)$ on MSE and cross-entropy losses.
[…]
It is unknown to the authors whether catapults occur for SignSGD, and if so, whether there is a “catapult triangle” in the phase diagram as seen with cross-entropy loss and SGD Figure 1(b,d).
Results of simulations of our one-parameter toy model with SignSGD are shown in Figure 11.
Figure 11: Simulations of our one-parameter model with SignSGD with MSE loss (left; see Equation 4) and cross-entropy loss (right; see Equation 9) with varying $\gamma,\eta$ exhibit predicted scaling relationships. Simulations run with $L=5$. Unlike with SGD, SignSGD does not diverge; instead, points in the upper region oscillate forever.
F.2 Empirical Plots of Phase Diagram Adam
Figure 12: Empirical runs of a 3 layer and readout ConvNet without residuals model trained with Adam with MSE loss (left; see Equation 4) and cross-entropy loss (right; see Equation 9) with varying $\gamma,\eta$ on CIFAR-5m.
We supplement the theoretical plots from Figure 11 with empirical plots in a Convolutional Neural Network with 3 layers and a readout layer, no residuals, trained on CIFAR-5m using Adam with MSE and Cross Entropy losses 12. The models are trained using learning rate warmup for approximately $5\%$ of the total number of steps, followed by a cosine decay until the end of training.
Appendix G Additional Results on Single-Index Models
In this section we empirically show that the results of the main text transfer to the commonly studied class of “single index models”. Here, we train a $3$-layer MLP of width 200 with on a single-index model of the form $y=(\bm{w}\cdot\bm{x})^{k}$ where $\bm{w}\sim\mathcal{N}(0,I)$, $\bm{x}\sim\mathcal{N}(0,I)$, $\bm{w}\in\mathbb{R}^{D}$, $x\in\mathbb{R}^{D}$. Our setting is the same as in the main text, namely the MLP is parameterized in $\mu$P and is centered. We show the phase portrait of training this model in an online setting with SGD in Figure 13. This portrait is consistent with that predicted by our linear model.
Figure 13: Empirical runs of a 3 layer and readout MLP model trained with SGD and MSE loss on a quadratic single-index model $y=(\bm{w}\cdot\bm{x})^{2}$. a) Verification of consistency across widths in $\mu$P b) The phase plot observed is as in Figure 1. c) and d) illustrate the same, but for $y=10^{-1}(\bm{w}\cdot\bm{x})^{3}$.
Appendix H Additional ResNet and ViT Plots